I checked out the elog from the vent in October 2016 when the OMC was removed from the path. In the vent in a couple weeks, we'd like to get the beam going through the OMC again. I wasn't really there for this last vent and don't have a great sense for how things go at the 40m, but this is how I think the procedure for this work should approximately go. The main points are that we'll need to slightly translate and rotate OM5, rotate OM6, replace one mirror that was removed last time, and add some beam dumps. Please let me know what I've got wrong or am missing.
[side note, I want to make some markup on the optics layouts that I see as pdfs elsewhere in the log and wiki, but haven't done it and didn't much want to dig around random drawing software, if there's a canonical way this is done please let me know.]
Steps to return the OMC to the IFO output: